Course: Acoustics 15. A circular membrane with radius 0,2 m; density 1,0 kg/m2; stretched to have a linear pressure of 10.000 N/m. What is the frequency of the second overtone mode for vibration circularly symmetric? a. 191,5 Hz b. 434 H:z c. 689 Hz d. 868 Hz e. 938,7 Hz
Answer #1

For second overtone, diameter = 5/2 x wavelength

wavelength = 2/5 diameter

v = /T

flambda = sqrt{T/ ho}

f= 1/lambda sqrt{T/ ho}

= 5/2d x /T

= 5/(2x0.4) x V/ 1 0000/ 1

= 625 Hz